find the sum of coefficient of binomial expansion of (1+x+x^2)^6

Grade:11

1 Answers

Tamer Eldebarky
6 years ago
The series (1+x+x2) represent geometric series with the first term a=1, common ratio r= x and the number of terms n=3. Its sum Sn=(a(1-rn))/(1-r)=(1)(1-x2)/(1-x)=((1-x)(1+x))/(1-x)=(1+x).
Then the sum of (1+x+x2)6=(1+x)6.